# PaCo-RL: Advancing Reinforcement Learning for Consistent Image Generation with Pairwise Reward Modeling

### v1

@@ 5.1 Experimental Setup
Models. We trained two variants of PaCo-Reward based on Qwen2.5-VL-7B-Instruct [bai2025qwen25vltechnicalreport]: (1) PaCo-Reward-7B-Fast, trained with only binary labels for fast convergence; (2) PaCo-Reward-7B, trained on the full PaCo-Dataset with reasoning-augmented labels for improved performance.
Benchmarks. We evaluated PaCo-Reward on two benchmarks: (1) ConsistencyRank, which contains $\sim$3k human-annotated instances, each containing one reference image and four candidate images ranked by visual consistency. (2) EditReward-Bench [wu2025editrewardhumanalignedrewardmodel], comprising 3k pairs of original and edited images with human preference labels over Prompt Following (PF) and Consistency (C).
Baselines. We compared PaCo-Reward against state-of-the-art reward models, including CLIP-I [radford2021learningtransferablevisualmodels], DreamSim [fu2023dreamsim], InternVL3.5-8B [wang2025internvl35advancingopensourcemultimodal], and Qwen2.5-VL-7B [bai2025qwen25vltechnicalreport] on ConsistencyRank; and GPT-4.1 [openai2024gpt4technicalreport], GPT-5 [OpenAI2025IntroducingGPT5], Gemini2.5-Pro [comanici2025gemini25pushingfrontier], the Qwen2.5-VL series [bai2025qwen25vltechnicalreport], and the EditScore series [luo2025editscoreunlockingonlinerl] on EditReward-Bench.
Metrics. We evaluate using Accuracy, Kendall’s rank correlation coefficient ($\tau$), Spearman’s rank correlation coefficient ($\rho$), and Top-1-Bottom-1 (T1-B1) Accuracy on ConsistencyRank, and PF, C, and Overall (geometric mean of PF and C) on EditReward-Bench, following [luo2025editscoreunlockingonlinerl].
@@
Models. For models, we used FLUX.1-dev [blackforestlabs_flux1dev_2024] for Text-to-ImageSet generation, and FLUX.1-Kontext-dev [labs2025flux1kontextflowmatching] as well as Qwen-Image-Edit [wu2025qwenimagetechnicalreport] for Image Editing tasks. PaCo-GRPO was employed to train these models using PaCo-Reward-7B as the reward model. The prompt templates for each task are provided in Appendix 8.
Benchmarks and Metrics. For benchmarks, we conducted evaluations on T2IS-Bench [jia2025settleonetexttoimagesetgeneration] for Text-to-ImageSet and GEdit-Bench [wu2025qwenimagetechnicalreport] for Image Editing. Evaluation metrics followed the corresponding benchmarks: comprehensive metrics from T2IS-Bench for quality and consistency, and Semantic Consistency (SC), Prompt Quality (PQ), Overall scores from GEdit-Bench for editing performance.
@@
SeedEdit3.0 [wang2025seededit30fasthighquality] 7.396 7.899 7.137 7.222 7.885 6.983 7.370 7.870 7.105 7.195 7.851 6.942
GPT-Image-1 [openai_gptimage1] 7.867 8.097 7.590 7.743 8.133 7.494 7.840 8.075 7.560 7.708 8.095 7.451
@@
OmniGen [xiao2024omnigenunifiedimagegeneration] 6.037 5.856 5.154 5.879 5.871 5.005 6.015 5.830 5.122 5.850 5.845 4.976
OmniGen2 [wu2025omnigen2explorationadvancedmultimodal] - - - 6.72 7.20 6.28 - - - - - -
OmniGen2 + EditScore [luo2025editscoreunlockingonlinerl] - - - 7.20 7.46 6.68 - - - - - -
Step1X-Edit [wu2025editrewardhumanalignedrewardmodel] 7.289 6.962 6.618 7.131 6.998 6.444 7.464 7.076 6.779 7.647 7.398 6.983
Step1X-Edit + EditReward [wu2025editrewardhumanalignedrewardmodel] 7.895 6.946 7.131 7.854 6.931 7.086 7.757 7.024 7.074 7.658 6.995 7.001
FLUX.1-Kontext [labs2025flux1kontextflowmatching] 6.842 6.888 6.143 6.599 6.873 5.956 - - - - - -
@@
Qwen-Image-Edit [wu2025qwenimagetechnicalreport] 7.746 7.805 7.307 7.642 7.807 7.223 7.727 7.977 7.344 7.633 7.938 7.264
@@ 5.2 Experimental Results
RA2: PaCo-Reward improves consistency in RL-based image generation. We integrate PaCo-Reward-7B into PaCo-GRPO for both Text-to-ImageSet generation and Image Editing tasks, as shown in Tabs.˜3 and 4. To ensure cross-model evaluation of visual consistency on T2IS-Bench, we additionally employ Gemma-3-4B-IT [gemmateam2025gemma3technicalreport], since PaCo-Reward-7B is fine-tuned from Qwen2.5-VL-7B.
On T2IS-Bench, FLUX.1-dev enhanced with PaCo-Reward-7B substantially outperforms the strongest open-source baseline [jia2025settleonetexttoimagesetgeneration] across all visual consistency metrics, achieving average absolute gains of 0.117 under the Qwen2.5-VL-7B evaluator and 0.103 under the Gemma-3-4B-IT evaluator. Moreover, it reaches performance comparable to closed-source models such as GPT-4, narrowing the gap between open-source and closed-source solutions.
On GEdit-Bench, PaCo-Reward-7B consistently improves SC and PQ across all base models. Unlike EditReward [wu2025editrewardhumanalignedrewardmodel], which degrades perceptual quality for Step1X-Edit, our method achieves a balanced improvement, enhancing consistency while preserving image quality. Even for strong baselines such as Qwen-Image-Edit, PaCo-Reward-7B further boosts both consistency and perceptual quality. These gains stem from the strong reward signal provided by PaCo-Reward-7B, which guides RL optimization toward consistent yet visually coherent outputs.
@@
Resolution-decoupled training: Fig.˜6 illustrates training under different image resolutions. Each training image uses a $2\times 2$ grid layout and is divided into four sub-images for reward computation, effectively halving the resolution in each dimension. As shown in Fig.˜6, low-resolution training ($512\times 512$) starts with lower rewards but reaches the performance of $1024\times 1024$ training after about 50 epochs, demonstrating that lower-quality images still provide reliable reward feedback [liu2025flowgrpotrainingflowmatching]. Low-resolution training also shows higher reward variance (indicated by circles in Figs.˜6 and 6), which promotes exploration and produces more diverse samples that enhance RL optimization. However, as shown in Fig.˜6, $256\times 256$ training fails due to insufficient visual detail for reliable reward evaluation. These results indicate that moderate resolution reduction can greatly accelerate training without degrading final performance.
Log-tamed multi-reward aggregation: To assess the effect of log-tamed aggregation in mitigating reward domination, we compare it with standard weighted-sum aggregation (see Fig.˜6). Two reward components are used: Consistency (PaCo-Reward-7B) and Prompt Alignment (CLIP-T [radford2021learningtransferablevisualmodels]). The plot shows the ratio of Consistency to Prompt Alignment rewards during training. Under standard aggregation without careful weight tuning, this ratio quickly exceeds 2.5 after 50 epochs, indicating that the Consistency reward dominates optimization. In contrast, log-tamed aggregation keeps the ratio below 1.8 throughout training, preventing any single reward from dominating and producing balanced improvements across all objectives.
@@ 5.3 Case Studies
Due to space constraints, we present three representative Text-to-ImageSet examples in Fig.˜7, with additional results for Text-to-ImageSet and Image Editing in Appendix LABEL:app:training_process_visualization. Fig.˜7 visualizes progressive improvements achieved during PaCo-RL training with a fixed random seed. In the identity case, the dentist’s facial and hairstyle attributes gradually converge toward a consistent appearance across different medical scenarios. In the style case, the chalkboard font across café menus becomes increasingly uniform, with higher prompt fidelity and improved overall aesthetics. In the logic case, the model learns a correct sketch-to-drawing progression, where each subsequent panel extends and refines the previous one rather than altering it. Together, these cases demonstrate PaCo-RL’s capability to jointly enhance diverse dimensions of visual consistency, validating effectiveness of both reward model and RL framework.
@@ 6 Conclusion
Our work introduces PaCo-RL, a novel reinforcement learning framework designed for consistent image generation. It consists of two key components: PaCo-Reward, a specialized reward model for pairwise visual consistency, and PaCo-GRPO, an online RL algorithm that enhances training efficiency and stability. Together, they enable more human-aligned and computationally efficient training for consistent image generation tasks. Extensive experiments demonstrate that our framework achieves state-of-the-art consistency performance with improved stability and efficiency, opening new avenues for advancing RL paradigms in consistent image generation.
Ref. Image
“Make him look stronger”
Ref. Image
“Make the action of the woman to laughing”
Ref. Image
“Make him gain 20 pounds”
Ref. Image
“Add abs to the original photo”
Figure 8: Training progression visualization for Image Editing with different prompts.
Training progression
Prompt: “Depict the chronological decomposition of a single leaf on a forest floor. All images maintain a realistic style with consistent lighting and environmental elements, focusing on the gradual transformation of the leaf while adhering to natural decay processes. The forest floor setting includes subtle elements like soil texture, scattered debris, and occasional fungi or insects.”
Figure 9: Training progression visualization for Text-to-ImageSet generation.
Training progression
Prompt: “Health beverage labels featuring honey drip font with viscous liquid texture and hexagonal comb patterns. All labels utilize the honey drip font style, integrating hexagonal comb motifs and natural/organic themes. Consistency in color palette (golden, amber, earthy tones) and texture emphasis ensures visual harmony across the set.”
Figure 10: Training progression visualization for Text-to-ImageSet generation.
Training progression
Prompt: “Step-by-step progression of creating a cheerful chef emoji. All images use a minimalist, cartoonish style with a clean white background. Bright and cohesive color schemes unify the stages, maintaining continuity in character proportions and playful energy.”
Figure 11: Training progression visualization for Text-to-ImageSet generation.
Prompt: “Please generate four different perspective images of a 3D animated parrot with a vibrant and colorful plumage. The parrot exhibits a stunning array of colors, including shades of red, green, blue, and yellow, with detailed feather textures that reflect light and give a sense of depth…”
Ours
AutoT2IS
Seedream
Figure 12: Comparison of different methods for multi-image generation.
Prompt: “Design product mockups featuring a retro, pixel art logo that reimagines our brand in an 8-bit style paired with a futuristic digital font. Apply the logo on 4 products: a portable gaming console, a vintage-style gaming t-shirt, a pixel art coffee mug, and a limited edition poster, using a monochromatic color scheme.”
Ours
AutoT2IS
Seedream
Figure 13: Comparison of different methods for multi-image generation.
Prompt: “This artwork represents the gradual creation of a traditional Chinese ink painting featuring pumpkins and vines. The progression follows these steps:…”
Ours
AutoT2IS
Seedream
Figure 14: Comparison of different methods for Text-to-ImageSet generation.
[…]
@@ 7 Implementation Details
PaCo-Reward Implementation Details. PaCo-Reward is fine-tuned using the LlamaFactory framework [zheng2024llamafactory] with a customized weighted cross-entropy loss defined in Sec.˜3.2. Both PaCo-Reward-7B-Fast and PaCo-Reward-7B are fine-tuned from Qwen2.5-VL-7B-Instruct [qwen2025qwen25technicalreport], employing a LoRA rank of 32, $\alpha=64$, a learning rate of $2\times 10^{-4}$, and a batch size of 8. Training on the PaCo-Dataset for two epochs requires approximately 18 GPU hours. For PaCo-Reward-7B, the first-token weight $\alpha$ in Sec.˜3.2 is set to 0.1 to emphasize the importance of the initial token. Since PaCo-Reward-7B-Fast is trained solely on binary labels (“Yes.”/“No.”), it does not apply token weighting.
PaCo-RL Implementation Details. During PaCo-RL training, one GPU is dedicated to running a vLLM [kwon2023efficient] server for reward computation, while the remaining seven GPUs are used for reinforcement-learning fine-tuning. Unless otherwise specified, all experiments adopt a learning rate of $3\times 10^{-4}$, a batch size of 1, a group size of $G=16$, and 42 unique samples per epoch. The clipping parameter $\varepsilon$ in Eq.˜1 is set to $1\times 10^{-4}$. Following [liu2025flowgrpotrainingflowmatching], we monitor KL loss throughout training but assign it zero weight ($\beta=0$) in Eq.˜1 to achieve better performance.
Text-to-ImageSet. For the Text-to-ImageSet generation task, we fine-tune FLUX.1-dev [blackforestlabs_flux1dev_2024] using LoRA with rank 64 and $\alpha=128$, and apply a classifier-free guidance (CFG) scale of 3.5 during both training and inference. The noise scale $\sigma_{t}$ in Sec.˜4.1 is defined as $\sigma_{t}=a\sqrt{\tfrac{t}{1-t}}$ with $a=0.7$, following FlowGRPO [liu2025flowgrpotrainingflowmatching]. We use 10 denoising steps and perform SDE sampling only at timestep $t=1$ (from $0,1,\ldots,9$}), leveraging MixGRPO [li2025mixgrpounlockingflowbasedgrpo] and FlowGRPO-Fast [liu2025flowgrpotrainingflowmatching] strategies for efficient training. In our PaCo-RL setup, the training resolution is $512\times 512$ while the evaluation resolution is $1024\times 1024$. We use $\delta=0.2$ in Eq.˜4 to aggregate the Consistency Score (from PaCo-Reward-7B) and the Text-Image Alignment Score (from CLIP-T [radford2021learningtransferablevisualmodels]).
Image Editing. For Image Editing, we fine-tune both FLUX.1-Kontext-dev [batifol2025flux] and Qwen-Image-Edit [wu2025qwenimagetechnicalreport] using LoRA with rank 64 and $\alpha=128$. The CFG scales are set to 2.5 and 4.0 for FLUX.1-Kontext-dev and Qwen-Image-Edit, respectively. We apply the MixGRPO [li2025mixgrpounlockingflowbasedgrpo] strategy for efficient training, setting the noise scale $a=0.9$ at timestep 1 for FLUX.1-Kontext-dev [batifol2025flux], and $a=1.0$ for timesteps 1-4 for Qwen-Image-Edit [wu2025qwenimagetechnicalreport]. The training and evaluation resolutions are $384\times 384$ and $1024\times 1024$, respectively. As the Image Editing task relies on a single reward signal, multi-reward aggregation and the log-tame strategy in Eq.˜4 are not employed. The prompt template for Image Editing reward computation is provided in Sec.˜8.
@@ 8 Prompt Templates
For the Image Editing task, we adopt a modified version of the prompt template from EditScore [luo2025editscoreunlockingonlinerl]. This template consists of two components: Semantic Consistency (SC) and Prompt Following (PF).

### v2

@@ 5.1 Experimental Setup
Models. We trained two variants of PaCo-Reward based on Qwen2.5-VL-7B-Instruct [1]: (1) PaCo-Reward-7B-Fast, trained with only binary labels for fast convergence; (2) PaCo-Reward-7B, trained on the full PaCo-Dataset with reasoning-augmented labels for improved performance.
Benchmarks. We evaluated PaCo-Reward on two benchmarks: (1) ConsistencyRank, which contains $\sim$3k human-annotated instances, each containing one reference image and four candidate images ranked by visual consistency. (2) EditReward-Bench [81], comprising 3k pairs of original and edited images with human preference labels over Prompt Following (PF) and Consistency (C).
Baselines. We compared PaCo-Reward against state-of-the-art reward models, including CLIP-I [52], DreamSim [15], InternVL3.5-8B [74], and Qwen2.5-VL-7B [1] on ConsistencyRank; and GPT-4.1 [44], GPT-5 [47], Gemini2.5-Pro [9], the Qwen2.5-VL series [1], and the EditScore series [40] on EditReward-Bench.
Metrics. We evaluate using Accuracy, Kendall’s rank correlation coefficient ($\tau$), Spearman’s rank correlation coefficient ($\rho$), and Top-1-Bottom-1 (T1-B1) Accuracy on ConsistencyRank, and PF, C, and Overall (geometric mean of PF and C) on EditReward-Bench, following [40].
@@
Models. For models, we used FLUX.1-dev [29] for Text-to-ImageSet generation, and FLUX.1-Kontext-dev [28] as well as Qwen-Image-Edit [78] for Image Editing tasks. PaCo-GRPO was employed to train these models using PaCo-Reward-7B as the reward model. The prompt templates for each task are provided in Appendix 8.
Benchmarks and Metrics. For benchmarks, we conducted evaluations on T2IS-Bench [24] for Text-to-ImageSet and GEdit-Bench [78] for Image Editing. Evaluation metrics followed the corresponding benchmarks: comprehensive metrics from T2IS-Bench for quality and consistency, and Semantic Consistency (SC), Prompt Quality (PQ), Overall scores from GEdit-Bench for editing performance.
@@
SeedEdit3.0 [73] 7.396 7.899 7.137 7.222 7.885 6.983 7.370 7.870 7.105 7.195 7.851 6.942
GPT-Image-1 [45] 7.867 8.097 7.590 7.743 8.133 7.494 7.840 8.075 7.560 7.708 8.095 7.451
@@
OmniGen [86] 6.037 5.856 5.154 5.879 5.871 5.005 6.015 5.830 5.122 5.850 5.845 4.976
OmniGen2 [79] - - - 6.72 7.20 6.28 - - - - - -
OmniGen2 + EditScore [40] - - - 7.20 7.46 6.68 - - - - - -
Step1X-Edit [81] 7.289 6.962 6.618 7.131 6.998 6.444 7.464 7.076 6.779 7.647 7.398 6.983
Step1X-Edit + EditReward [81] 7.895 6.946 7.131 7.854 6.931 7.086 7.757 7.024 7.074 7.658 6.995 7.001
FLUX.1-Kontext [28] 6.842 6.888 6.143 6.599 6.873 5.956 - - - - - -
@@
Qwen-Image-Edit [78] 7.746 7.805 7.307 7.642 7.807 7.223 7.727 7.977 7.344 7.633 7.938 7.264
@@ 5.2 Experimental Results
RA2: PaCo-Reward improves consistency in RL-based image generation. We integrate PaCo-Reward-7B into PaCo-GRPO for both Text-to-ImageSet generation and Image Editing tasks, as shown in Tabs.˜3 and 4. To ensure cross-model evaluation of visual consistency on T2IS-Bench, we additionally employ Gemma-3-4B-IT [64], since PaCo-Reward-7B is fine-tuned from Qwen2.5-VL-7B.
On T2IS-Bench, FLUX.1-dev enhanced with PaCo-Reward-7B substantially outperforms the strongest open-source baseline [24] across all visual consistency metrics, achieving average absolute gains of 0.117 under the Qwen2.5-VL-7B evaluator and 0.103 under the Gemma-3-4B-IT evaluator. Moreover, it reaches performance comparable to closed-source models such as GPT-4, narrowing the gap between open-source and closed-source solutions.
On GEdit-Bench, PaCo-Reward-7B consistently improves SC and PQ across all base models. Unlike EditReward [81], which degrades perceptual quality for Step1X-Edit, our method achieves a balanced improvement, enhancing consistency while preserving image quality. Even for strong baselines such as Qwen-Image-Edit, PaCo-Reward-7B further boosts both consistency and perceptual quality. These gains stem from the strong reward signal provided by PaCo-Reward-7B, which guides RL optimization toward consistent yet visually coherent outputs.
@@
Resolution-decoupled training: Fig.˜6 illustrates training under different image resolutions. Each training image uses a $2\times 2$ grid layout and is divided into four sub-images for reward computation, effectively halving the resolution in each dimension. As shown in Fig.˜6, low-resolution training ($512\times 512$) starts with lower rewards but reaches the performance of $1024\times 1024$ training after about 50 epochs, demonstrating that lower-quality images still provide reliable reward feedback [37]. Low-resolution training also shows higher reward variance (indicated by circles in Figs.˜6 and 6), which promotes exploration and produces more diverse samples that enhance RL optimization. However, as shown in Fig.˜6, $256\times 256$ training fails due to insufficient visual detail for reliable reward evaluation. These results indicate that moderate resolution reduction can greatly accelerate training without degrading final performance.
Log-tamed multi-reward aggregation: To assess the effect of log-tamed aggregation in mitigating reward domination, we compare it with standard weighted-sum aggregation (see Fig.˜6). Two reward components are used: Consistency (PaCo-Reward-7B) and Prompt Alignment (CLIP-T [52]). The plot shows the ratio of Consistency to Prompt Alignment rewards during training. Under standard aggregation without careful weight tuning, this ratio quickly exceeds 2.5 after 50 epochs, indicating that the Consistency reward dominates optimization. In contrast, log-tamed aggregation keeps the ratio below 1.8 throughout training, preventing any single reward from dominating and producing balanced improvements across all objectives.
@@ 5.3 Case Studies
Fig.˜7 visualizes progressive improvements achieved during PaCo-RL training with a fixed random seed. In the identity case, the dentist’s facial and hairstyle attributes gradually converge toward a consistent appearance across different medical scenarios. In the style case, the chalkboard font across café menus becomes increasingly uniform, with higher prompt fidelity and improved overall aesthetics. In the logic case, the model learns a correct sketch-to-drawing progression, where each subsequent panel extends and refines the previous one rather than altering it. Together, these cases demonstrate PaCo-RL’s capability to jointly enhance diverse dimensions of visual consistency, validating effectiveness of both reward model and RL framework.
@@ 6 Conclusion
Our work introduces PaCo-RL, a novel reinforcement learning framework designed for consistent image generation. It consists of PaCo-Reward and PaCo-GRPO, which together enable more human-aligned and computationally efficient training for consistent image generation tasks.
Acknowledgment
This work is supported by the Fundamental and Interdisciplinary Disciplines Breakthrough Plan of the Ministry of Education of China (No. JYB2025XDXM101), the National Natural Science Foundation of China (No. 62272374, No. 62192781), the Natural Science Foundation of Shaanxi Province (No.2024JC-JCQN-62), the State Key Laboratory of Communication Content Cognition under Grant No. A202502, the Key Research and Development Project in Shaanxi Province (No. 2023GXLH-024) and by A*STAR Career Development Fund <Project No. C243512010>.
@@ 7 Implementation Details
PaCo-Reward Implementation Details. PaCo-Reward is fine-tuned using the LlamaFactory framework [96] with a customized weighted cross-entropy loss defined in Sec.˜3.2. Both PaCo-Reward-7B-Fast and PaCo-Reward-7B are fine-tuned from Qwen2.5-VL-7B-Instruct [51], employing a LoRA rank of 32, $\alpha=64$, a learning rate of $2\times 10^{-4}$, and a batch size of 8. Training on the PaCo-Dataset for two epochs requires approximately 18 GPU hours. For PaCo-Reward-7B, the first-token weight $\alpha$ in Sec.˜3.2 is set to 0.1 to emphasize the importance of the initial token. Since PaCo-Reward-7B-Fast is trained solely on binary labels (“Yes.”/“No.”), it does not apply token weighting.
PaCo-RL Implementation Details. During PaCo-RL training, one GPU is dedicated to running a vLLM [27] server for reward computation, while the remaining seven GPUs are used for reinforcement-learning fine-tuning. Unless otherwise specified, all experiments adopt a learning rate of $3\times 10^{-4}$, a batch size of 1, a group size of $G=16$, and 42 unique samples per epoch. The clipping parameter $\varepsilon$ in Eq.˜1 is set to $1\times 10^{-4}$. Following [37], we monitor KL loss throughout training but assign it zero weight ($\beta=0$) in Eq.˜1 to achieve better performance.
Text-to-ImageSet. For the Text-to-ImageSet generation task, we fine-tune FLUX.1-dev [29] using LoRA with rank 64 and $\alpha=128$, and apply a classifier-free guidance (CFG) scale of 3.5 during both training and inference. The noise scale $\sigma_{t}$ in Sec.˜4.1 is defined as $\sigma_{t}=a\sqrt{\tfrac{t}{1-t}}$ with $a=0.7$, following FlowGRPO [37]. We use 10 denoising steps and perform SDE sampling only at timestep $t=1$ (from $0,1,\ldots,9$}), leveraging MixGRPO [31] and FlowGRPO-Fast [37] strategies for efficient training. In our PaCo-RL setup, the training resolution is $512\times 512$ while the evaluation resolution is $1024\times 1024$. We use $\delta=0.2$ in Eq.˜4 to aggregate the Consistency Score (from PaCo-Reward-7B) and the Text-Image Alignment Score (from CLIP-T [52]).
Image Editing. For Image Editing, we fine-tune both FLUX.1-Kontext-dev [2] and Qwen-Image-Edit [78] using LoRA with rank 64 and $\alpha=128$. The CFG scales are set to 2.5 and 4.0 for FLUX.1-Kontext-dev and Qwen-Image-Edit, respectively. We apply the MixGRPO [31] strategy for efficient training, setting the noise scale $a=0.9$ at timestep 1 for FLUX.1-Kontext-dev [2], and $a=1.0$ for timesteps 1-4 for Qwen-Image-Edit [78]. The training and evaluation resolutions are $384\times 384$ and $1024\times 1024$, respectively. As the Image Editing task relies on a single reward signal, multi-reward aggregation and the log-tame strategy in Eq.˜4 are not employed. The prompt template for Image Editing reward computation is provided in Sec.˜8.
@@ 8 Prompt Templates
For the Image Editing task, we adopt a modified version of the prompt template from EditScore [40]. This template consists of two components: Semantic Consistency (SC) and Prompt Following (PF).
@@
9 Resolution-Decoupled Training Analysis
To validate the reliability of resolution-decoupled training, we conduct experiments generating image sets at three resolutions: $256\times 256$ (0.25x), $512\times 512$ (0.5x), and $1024\times 1024$ (1x) using FLUX.1-dev, and analyze the Pearson correlation of evaluation metrics across resolutions.
(a) 1x vs. 0.5x
(b) 1x vs. 0.25x
Figure 15: Pearson correlation of evaluation metrics across different training-to-inference resolution ratios. Strong correlations at 0.5x confirm that reward signals remain reliable under moderate resolution reduction.
As shown in Fig.˜15, all metrics exhibit strong positive correlations (0.7250.7250.8480.848) when comparing 1x vs. 0.5x resolutions, indicating that the relative quality ordering of image sets is largely preserved and reward signals remain reliable under moderate resolution reduction. However, correlations weaken substantially when comparing 1x vs. 0.25x, especially for Aesthetics and Consistency, suggesting that extremely low resolutions lose detailed visual information crucial for these fine-grained aspects.
We further validated these findings on additional reward models (Text-Rendering, GenEval) and generators (Qwen-Image, FLUX2), confirming that 0.5x reduction consistently preserves reward reliability while 0.25x fails, particularly for fine-grained metrics. Regarding higher-resolution inference, no systematic artifacts were observed when generating at full resolution after 0.5x training, as the reward improvements transfer effectively across resolutions. This supports the use of resolution-decoupled training as a practical strategy for reducing computational cost without sacrificing final generation quality.
10 Ablation on PaCo-GRPO Components
We quantify the individual contributions of two key components in PaCo-GRPO: resolution-decoupled training (Res-Dec.) and log-tamed reward aggregation (Log-Agg.). Results are reported in Tab.˜5.
Table 5: Ablation study on PaCo-GRPO components. Removing resolution-decoupled training leads to suboptimal performance even with doubled training time, while removing log-tamed aggregation causes reward collapse.
[…]
Full PaCo-GRPO 0.555 0.728 0.493 6.0h
W/O Res-Dec. 0.542 0.698 0.452 12.0h
W/O Log-Agg. 0.471 0.616 0.557 6.0h
Resolution-Decoupled Training.
Removing resolution-decoupled training results in suboptimal performance even with doubled training time (12h vs. 6h), as the model fails to converge fully at higher resolution. This demonstrates that training at a reduced resolution (0.5x) not only halves the computational cost but also leads to better optimization dynamics by enabling more effective exploration within the same time budget.
Log-Tamed Reward Aggregation.
Without log-tamed aggregation, the visual consistency reward dominates the training signal, causing the model to collapse into generating near-identical, low-quality image sets. While the Visual Consistency (V.C.) score increases to 0.557, both Aesthetics and Prompt Following degrade significantly, confirming that the log-tamed strategy in Eq.˜4 is essential for balancing multiple reward objectives and preventing reward hacking.